if 180<theta<270, cot theta=4/3 then find the value of sin theta/3
a.3/√10
b.1/√5

Grade:11

1 Answers

Aditya Gupta
5 years ago
cot theta=4/3
cosec^2theta=cot^2theta+1=25/9
so sintheta= – 3/5
now, sintheta= 3sin theta/3 – 4sin^3 theta/3= – 3/5
or sin theta/3= 0.95261